OpenStudy (anonymous):

What is the product in simplest form? State any restrictions on the variable. \frac{ y^2 }{ y-3 } times \frac{ y^2 - y - 6 }{ y^2 + 1y }

OpenStudy (anonymous):

\[ \frac{ y^2 }{ y-3 } \times \frac{ y^2 - y - 6 }{ y^2 + 1y }\]

ganeshie8 (ganeshie8):

factor numerator and denominator... on right side fraction

ganeshie8 (ganeshie8):

factor : \(\large y^2-y-6\) find two numbers such that, they multiply to \(-6\) they add up to \(-1\)

OpenStudy (whpalmer4):

Before you do any cancellation (assuming there are factors that cancel), take a look at your factored denominators and check to see what values of \(y\) (if any) will cause the denominator to equal 0. Those are your restricted variables, and they apply even for factors that may be canceled out of the final, simplest form. It's easy to forget to check the ones that aren't present at the end!

OpenStudy (anonymous):

\[\frac{ y^2 }{ y - 3 } \times \frac{ (y + 2 )(y - 3) }{(y+1)(y-1) }\] I think the numerator is correct but for some reason I'm stuck on the denominator because I think it's supposed to be two numbers that equal a positive one when you add and multiply it... I'm not sure >.<

OpenStudy (anonymous):

Oh wait i didn't you that you typed lol

OpenStudy (anonymous):

No... still confused /.\

ganeshie8 (ganeshie8):

yes numerator is correct. denominator is just : \(y^2 + 1y = y(y+1)\)

ganeshie8 (ganeshie8):

and as whpalmer is saying, before cancelling itself, u need to check the 'restricted values'

OpenStudy (anonymous):

Oooooh I always forget that that's an option >-\ So... \[\frac{ y^2 }{ y-3 } \times \frac{ (y + 2)(y-3) }{ y(y + 1) }\] Who's Whpalmer? o. o And I'm sorry but how do I locate restricted values as well as... cancel them out?

ganeshie8 (ganeshie8):

nope. cancelling is a different part of the story

ganeshie8 (ganeshie8):

to find restricted values of a fraction : look at denominator

ganeshie8 (ganeshie8):

wat values of y will make the denominator become 0 ?

ganeshie8 (ganeshie8):

if u put y = 3, denominator becomes 0 right ?

ganeshie8 (ganeshie8):

if u put y = -1, denominator becomes 0 right ?

ganeshie8 (ganeshie8):

if u put y = 0, denominator becomes 0 rihgt ?

ganeshie8 (ganeshie8):

so, \( 3, -1, 0\) are the 'restricted values' for variable \(y\)

ganeshie8 (ganeshie8):

if that makes any sense at all...

ganeshie8 (ganeshie8):

they're called restricted values cuz, having \(0\) in the denominator is illegal in math.

ganeshie8 (ganeshie8):

basically, to find restricted values, u wud simply set the denominator to 0, and solve for the variable.

OpenStudy (anonymous):

Why does it have to be a 3 or -1? And oops I see Whpalmer now... my fault. Sorries... I didn't read the name. But... the denominator on both sides or just one?

ganeshie8 (ganeshie8):

\(\large \frac{ y^2 }{\color{green}{ y-3} } \times \frac{ (y + 2)(y-3) }{ \color{green}{y(y + 1)} } \)

ganeshie8 (ganeshie8):

that whole thing is the denominator, so u need to set the whole thing to 0, and solve for y

ganeshie8 (ganeshie8):

\(\large (y-3) y(y+1) = 0 \)

ganeshie8 (ganeshie8):

solve y